# HCOA*: Hierarchical Class-ordered A* for Navigation in Semantic Environments

### v1

@@ V Semantic Class Prediction
Let $\phi_{V}^{\ell}:V^{\ell}\to\mathcal{K}$ be the function that assigns semantic classes to nodes in the layers $\ell\neq 0$. Ideally, given an optimal path $\pi^{*}(v_{s},v_{g})$ computed in $G^{0}$, the semantic class of a node $P\in V^{\ell}$ is determined conservatively by selecting the highest semantic class present in the nodes in $\ell=0$ that both belong to the optimal path and are part of the subgraph $G^{0^{\prime}}(P)=\{u\in V^{0}:p(u,\ell)=P\}$ induced by $P$ in layer $\ell=0$. Formally, this is expressed as:
@@
$\pi^{\prime}(P)=\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}.$ (3b)
This process follows a bottom-up approach, necessitating the computation of the optimal path in layer $\ell=0$. Given that HCOA* works top-down, we seek an alternative method to assign semantic classes to the higher layers. In the next sections, we present the proposed approaches to predict the semantic classes of higher-layer nodes.
V-A Graph Neural Network Method
We formulate this problem as a supervised node classification task and design a GNN to predict the semantic classes. We construct a dataset $\mathcal{D}=\{G^{0^{\prime}}(P),\phi_{V}^{\ell}(P)\}$ by selecting nodes $P\in V^{\ell}$ for $\ell\neq 0$ and extracting their induced subgraphs $G^{0^{\prime}}(P)$. We then assign semantic classes to the nodes in $G^{0^{\prime}}(P)$ based on the given task. For instance, if the task is navigation with object avoidance, we can assign different priorities to regions by placing disks of a specified radius and semantic class around certain object nodes. Next, we execute COA* on this graph and compute $\phi_{V}^{\ell}(P)$, where the starting location is a border node.
A border node $v^{*}\in V^{0}$ of layer $\ell$ is defined as a node that shares an edge with another node from a different ancestor. Formally, there exists $e=(v^{*},u^{*})$ where $v^{*},u^{*}\in V^{0}$ and $p(v^{*},\ell)\neq p(u^{*},\ell)$. Border nodes play a crucial role in the classification of $P\in V^{\ell}$. If the start and goal nodes do not share the same ancestor, the optimal path must traverse a border node in layer $\ell$. Conversely, if both nodes have $P$ as their ancestor, the classification of $P$ becomes irrelevant, as HCOA* will not need to compute a path in layer $\ell$.
The proposed model is illustrated in Figure 3. We use border nodes as input features for the network, which are concatenated with the semantic classes of the nodes in $G^{0^{\prime}}$ represented as one-hot encodings. A 2-layer MLP is used to preprocess the input.
The GNN consists of three layers, each incorporating a Graph Convolutional Network (GCN) operator [17] for message passing, followed by batch normalization and a ReLU activation. To mitigate oversmoothing, we introduce skip connections between the GNN layers. Additionally, we apply average pooling to handle environments of varying sizes. The pooled representation is then processed through another 2-layer MLP, followed by a softmax function. Finally, we use cross-entropy as the loss function for node classification. Formally, given a dataset $\mathcal{D}$ and learnable parameters $\vartheta_{\text{GNN}},\vartheta_{\text{MLP}}$, the predicted semantic class is obtained by:
where $\widehat{\phi}_{V}^{\ell}=f(G^{0^{\prime}}(P);\vartheta_{\text{GNN}},\vartheta_{\text{MLP}})$ represents the model function parametrized by $\vartheta_{\text{GNN}},\vartheta_{\text{MLP}}$.
V-B Majority-Class Method
The Majority-Class (MC) method is a simple and fast approach for predicting the semantic class of a node $P\in V^{\ell}$ for $\ell\neq 0$ by directly calculating the semantic classes of the nodes in $G^{0^{\prime}}(P)$. Specifically, the predicted class is determined by counting the occurrences of each semantic class among the nodes in the induced subgraph $G^{0^{\prime}}(P)$ and selecting the most frequent one. That is,
@@ VI Simulations
We perform two sets of simulations on the publicly available uHumans2 office 3DSG, shown in Figure 1 and constructed using Hydra [2]. In the figure, the corresponding layers are also depicted. The places layer $\ell=0$ is a subgraph where each node represents an obstacle-free location, and edges indicate straight-line traversability. The rooms at layer $\ell=1$ consist of nodes representing room centers, with edges connecting neighboring rooms. The number of border nodes (bn) and places in each room are: R(0): 49 bn, 330 places, R(1): 22 bn, 259 places, R(2): 30 bn, 196 places, R(3): 32 bn, 195 places, R(4): 4 bn, 63 places, R(5): 35 bn, 206 places, R(6): 14 bn, 65 places.
We consider a set of three semantic classes $\mathcal{K}$, ordered by decreasing priority: 1 (Green), 2 (Blue), 3 (Red). We utilize the objects layer to identify objects and assign semantic classes to the surrounding place nodes.
In the first section, we present the training results of the GNN and its comparison with the MC approach. The second section demonstrates a path-planning scenario on the entire 3DSG and compares HCOA*, using both the GNN and MC methods, against COA*. HCOA*-GNN utilizes the best GNN model from the previous training phase for room inference. All the simulations were performed using Python 3.8.10 and PyTorch 2.4.1 on a computer with 2.2 GHz, 12-Core, Intel Core i7-8750H CPU, 16GB RAM and an Nvidia RTX 2060 GPU, 6GB VRAM.
@@ VI-A Performance Metrics
To evaluate the performance of room classification, we compute the accuracy, which measures the proportion of correctly classified rooms in each dataset:
$\text{Accuracy}=\frac{\sum\limits_{i=1}^{n}\mathbbm{1}(\widehat{\phi}_{V}^{\ell}(P_{i})=\phi_{V}^{\ell}(P_{i}))}{n},$ (6)
where $n$ is the number of samples, and $\mathbbm{1}(\cdot)$ is the indicator function, which equals 1 if the predicted class matches the ground truth and 0 otherwise.
For the path-planning scenario in the whole 3DSG, we compute the computational time of the algorithms along with the number of expanded nodes (Line 8 in Algorithm 1).
VI-B Room Node Classification
TABLE I: Room Node Classification.
Metrics GNN MC
Training Time (min) 90 -
Validation Accuracy (%) 67.64 30.36
Test Accuracy (Rooms with 1-20 bn) (%) 82.50 54.25
Test Accuracy (Rooms with 21-30 bn) (%) 65.00 19.75
Test Accuracy (Rooms with 31-40 bn) (%) 64.00 19.25
Test Accuracy (Rooms with 41-50 bn) (%) 59.00 24.50
In this section, we present the results from predicting the semantic class of room nodes. To generate the dataset, we constructed 14,000 graphs by extracting the induced subgraphs from all rooms and randomly assigning semantic classes to nodes within a disk of randomly chosen center locations, repeating this process 2,000 times per room. Then, we ran COA* to determine the semantic class of the room, starting from a randomly selected border node. The dataset was split into 80% training, 10% validation, and 10% testing. The model was trained using the Adam optimizer [18] with a learning rate of $10^{-2}$, over 1,600 epochs, a dropout rate of 0.2 and a batch size of 64. The GNN and MLP layers contain 32 neurons.
Table I summarizes the results of room node classification. For the test set, we analyze performance separately for rooms with different numbers of border nodes. The results indicate that the GNN approach outperforms significantly the MC method across both the validation and test sets. Additionally, the training accuracy of the GNN is 67.94%.
@@ VI-C Path-Planning on uHumans2 Office Scene
Figure 4: (a) Places subgraph of the uHumans2 office scene (52m×45m) with the assigned semantic classes (green, blue, red). The starting node is denoted by S, and the goal node by G; (b-d) The path of each algorithm is shown in bold black. Expanded nodes are depicted in regular black, while unexpanded nodes in gray: (b) Path of COA* on the places subgraph; (c) Path of HCOA* on the rooms subgraph; (d) Path of HCOA* on the places subgraph.
@@
The computed path of COA* in the places layer is shown in Figure 4(4(b)). Figures 4(4(c))-(4(d)) illustrate the results of HCOA* in the rooms and places layers, respectively. The figures also show the expanded nodes for each algorithm. In particular, COA * expanded nodes in two additional rooms compared to HCOA*. The exact number of expanded nodes is provided in Table II, where HCOA* demonstrates a 25% reduction in expanded nodes compared to COA*.
@@
Table II also presents the computational time of the algorithms over 1,000 runs. We observe that HCOA*-MC achieves the best performance in terms of computational efficiency, reducing the execution time by 16% compared to COA*. However, HCOA*-GNN has the highest computational time, as GNN inference can be more time-consuming than graph search in small graphs (the entire places subgraph contains only 1,314 nodes). It should be noted that the first run of HCOA*-GNN takes longer, likely due to compilation and GPU/CPU warm-up overhead. Additionally, both GNN and MC classify R(0) as class 2 and R(2) as class 3. However, while the GNN assigns class 3 to R(3), MC classifies it as class 1. The remaining rooms are not classified, as they are not adjacent to any expanded nodes.
@@
Finally, all three approaches successfully compute the optimal path in the places layer, as shown in Figure 4. It is important to note that while the first assumption of Proposition 2 holds, the second assumption does not. However, this does not affect the results in this case, as the optimal path in the places layer does not require deviations into rooms outside the unique path in the rooms layer to achieve a lower-cost trajectory.
TABLE II: Path-Planning on uHumans2 Office Scene.
Metrics HCOA*-GNN HCOA*-MC COA*
Expanded Nodes 412 412 549
Time ($10^{-3}$s) 38.8 $\pm$ 47.1 4.3 $\boldsymbol{\pm}$ 2.6 5.1 $\pm$ 2.5
Optimal Path
VII Conclusions

### v2

@@ V Semantic Class Prediction
Let $\phi_{V}^{\ell}:V^{\ell}\to\mathcal{K}$ be the function that assigns semantic classes to nodes in layers $\ell=1,\ldots,n$. Ideally, given an optimal path $\pi^{*}(v_{s},v_{g})$ computed in $G^{0}$, the semantic class of a node $P\in V^{\ell}$ ($\ell=1,\ldots,n$) is determined conservatively by selecting the highest semantic class present in the nodes in $\ell=0$ that both belong to the optimal path and are part of the subgraph $G^{0^{\prime}}(P)=\{u\in V^{0}:p(u,\ell)=P,\>\text{where}\>P\in V^{\ell}\}$ induced by $P$ on layer 0. Formally,
@@
$\pi^{\prime}(P)=\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}(P).$ (3b)
This process follows a bottom-up approach, necessitating the computation of the optimal path in layer $\ell=0$. Given that HCOA* works top-down, we seek alternative methods to find the semantic classes of the higher layers. We introduce three methods for predicting the semantic class of higher-layer nodes: a Majority-Class (MC), a k-Nearest Neighbors (kNN), and a Graph Neural Network (GNN) approach. The latter two methods require a supervised training phase, which necessitates the construction of a dataset.
To this end, we construct a dataset $\mathcal{D}=\{G^{0^{\prime}}(P),\phi_{V}^{\ell}(P)\}$ by selecting nodes $P\in V^{\ell}$ for $\ell\neq 0$ and extracting their induced subgraphs $G^{0^{\prime}}(P)$. We then assign semantic classes to the nodes in $G^{0^{\prime}}(P)$ based on the given task. For navigation tasks involving object avoidance, we determine region priorities by placing disks of a specified radius and semantic class around certain objects. Next, we run COA* on $G^{0^{\prime}}(P)$ and compute $\phi_{V}^{\ell}(P)$, where the start is a border node.
A border node $v^{*}\in V^{0}$ is a node that is connected by an edge to another (border) node $u^{*}\in V^{0}$ whose ancestor in layer $\ell$ is different from that of $v^{*}$. Formally, there exists $e=(v^{*},u^{*})$ where $v^{*},u^{*}\in V^{0}$ and $p(v^{*},\ell)\neq p(u^{*},\ell)$. Border nodes play a crucial role in the classification of $P\in V^{\ell}$. If the start and goal nodes do not share the same ancestor, the optimal path must traverse a border node in layer $\ell$. Conversely, if both nodes have $P$ as their ancestor, the classification of $P$ becomes irrelevant, as HCOA* will not need to compute a path in layer $\ell$.
V-A Majority-Class Method
The Majority-Class (MC) method is a simple and fast approach for predicting the semantic class of a node $P\in V^{\ell}$ for $\ell\neq 0$. The predicted class is determined by counting the occurrences of each semantic class in the induced subgraph $G^{0^{\prime}}(P)$ and selecting the most frequent one. That is,
@@ V-A Majority-Class Method
V-B k-Nearest Neighbors Method
The kNN algorithm [19] is a non-parametric, instance-based learning method that classifies data points based on the majority vote of their $\mathsf{k}$ nearest neighbors in the feature space. Given the induced subgraph $G^{0^{\prime}}(P)$ with node set $V^{0^{\prime}}$, we extract a graph-level feature vector $\mathsf{f}$ consisting of: (i) the proportion of each semantic class in the graph
$\frac{N_{V}(G^{0^{\prime}}(P),k)}{|V^{0^{\prime}}|},\quad k\in\mathcal{K},$ (5)
and (ii) the majority class among the border nodes $\operatorname*{arg\,max}_{k\in\mathcal{K}}{N_{V^{*}}(G^{0^{\prime}}(P),k)}$, where $V^{*}$ denotes the set of border nodes and $k$ the index of a semantic class (not to be confused with $\mathsf{k}$ used in kNN). We intentionally limit the number of features, as the inference time of kNN increases with the dimensionality of the feature space.
To ensure uniform scaling, we standardize all feature vectors $\mathsf{f}$. Graph similarity is measured via Euclidean distance, and classification is performed by assigning the majority class among the $\mathsf{k}$ nearest neighbors from the training set:
$\widehat{\phi}_{V}^{\ell}(P)=\operatorname*{arg\,max}\limits_{k\in\mathcal{K}}\sum\limits_{i=1}^{\mathsf{k}}\mathbbm{1}[\phi_{V}^{\ell}(P_{i})=k],$ (6)
where $P_{1},P_{2},\ldots,P_{\mathsf{k}}$ are the $\mathsf{k}$ nearest neighbors of $P$ in the training set, and $\mathbbm{1}[\cdot]$ is the indicator function that equals 1 if the condition is true and 0 otherwise.
V-C Graph Neural Network Method
The GNN method is a supervised learning approach that leverages the graph structure and features through message passing. The proposed model is shown in Figure 4. We use border nodes as input features, which are concatenated with the semantic classes of the nodes in $G^{0^{\prime}}$ represented as one-hot encodings. A 2-layer MLP is used to preprocess the input.
The GNN consists of three layers, each incorporating a Graph Convolutional Network (GCN) operator [20] for message passing, followed by batch normalization and a ReLU activation. We adopt a 3-layer architecture, as deeper GNNs tend to suffer from oversmoothing and slower inference, both of which are critical limitations in path-planning applications. To mitigate oversmoothing, we introduce skip connections between the GNN layers. Additionally, we apply average pooling to handle environments of varying sizes. The pooled representation is then processed through another 2-layer MLP, followed by a softmax function. Finally, we use cross-entropy as the loss function for node classification. Formally, given a dataset $\mathcal{D}$ and learnable parameters $\vartheta_{\text{GNN}},\vartheta_{\text{MLP}}$, the predicted semantic class is obtained by
where $\widehat{\phi}_{V}^{\ell}(P)=f(G^{0^{\prime}}(P);\vartheta_{\text{GNN}},\vartheta_{\text{MLP}})$ represents the model function parametrized by $\vartheta_{\text{GNN}},\vartheta_{\text{MLP}}$.
@@ VI Simulations
We perform simulations on two publicly available datasets generated with Hydra [2]: the uHumans2 office 3DSG (Figure 1) and the subway 3DSG (Figure 6(6(a))). Figure 1 also shows the layer names. The place layer ($\ell=0$) is a graph with nodes as obstacle-free locations and edges indicating traversability. The room layer ($\ell=1$) consists of nodes representing room centers and edges connecting neighboring rooms.
Our task involves navigation with object or room avoidance. The pipeline proceeds as follows: we first translate safety-related commands (e.g., “avoid computers”) into semantic classes at the place layer, and impose a total order based on their priority. Next, depending on the room classification method employed, we train the corresponding models to predict semantic room classes. Finally, we select the best-performing model and execute HCOA* to compute the path.
In all simulations, we impose two safety constraints resulting in three semantic classes $\mathcal{K}$, ordered by decreasing priority. We leverage the room and object layers to identify relevant rooms and objects, and assign semantic labels to nearby place nodes accordingly. The resulting classes are visualized (in decreasing priority) as: 1 (Green), 2 (Blue), and 3 (Red).
The first section presents the comparison of the three methods for semantic class prediction. The second section showcases path-planning scenarios in the office 3DSG, comparing HCOA* to COA*. We use the same weight function across all graph layers and denote it by $w$ for simplicity, corresponding to the Euclidean distance between nodes. We also include results from a modified A* algorithm, denoted MA* (similar to [9]), to demonstrate the importance of imposing a total order over semantic classes. In this variant, the edge cost in A* is modified to $w^{\prime}=w+\alpha^{k}$, where $k\in\mathcal{K}$ and $\alpha$ is a scaling factor. The third section presents analogous experiments in the subway 3DSG, also comparing HCOA* to COA* and MA*. In these experiments, HCOA*-GNN utilizes the best GNN model from the previous section for room inference.
All the simulations were performed using Python 3.8.10 and PyTorch 2.4.1 on a computer with 2.2 GHz, 12-Core, Intel Core i7-8750H CPU, 16GB RAM and an Nvidia RTX 2060 GPU, 6GB VRAM.
@@ VI-A Performance Metrics
To evaluate the performance of the prediction of a room’s semantic class, we compute the accuracy, which measures the proportion of correctly classified rooms in each dataset:
$\text{Accuracy}=\frac{\sum\limits_{i=1}^{n}\mathbbm{1}[\widehat{\phi}_{V}^{\ell}(P_{i})=\phi_{V}^{\ell}(P_{i})]}{n},$ (8)
where $n$ is the number of samples. For the planning scenarios, we compute the algorithms’ computational time along with the number of expanded nodes (Line 8 in Algorithm 1).
VI-B Semantic Class Prediction
TABLE I: Semantic Class Prediction.
Metrics MC kNN GNN
Training Time - 10 s 97 min
Validation Acc. (%) 42.79 52.43 63.43
Test Acc. (1-20 bn Rooms) (%) 57.00 70.50 74.00
Test Acc. (21-30 bn Rooms) (%) 33.75 48.50 58.25
Test Acc. (31-40 bn Rooms) (%) 37.00 44.25 54.75
Test Acc. (41-50 bn Rooms) (%) 41.50 43.00 47.50
In this section, we present the results from predicting the semantic class of room nodes. To generate the dataset, we constructed 14,000 graphs by extracting the induced subgraphs from all rooms in the Office 3DSG and randomly assigning semantic classes to nodes within a disk of randomly chosen center locations, repeating this process 2,000 times per room. Then, we ran COA* to determine the semantic class of the room, starting from a randomly selected border node. The dataset was split into 80% training, 10% validation, and 10% testing. The GNN model was trained using the Adam optimizer [21] with a learning rate of $10^{-2}$, over 1,600 epochs, a dropout rate of 0.2 and a batch size of 64. The GNN and MLP layers contain 32 neurons. For training the kNN model, we set $\mathsf{k}=5$, and used half of the training dataset.
Table I summarizes the results of room node classification. For the test set, we analyzed the performance separately for rooms with different numbers of border nodes (bn). The results indicate that the GNN approach outperforms the other two methods. This result is expected, as the GNN learns complex spatial, structural, and semantic features, unlike the MC baseline and the kNN method, which relies on hand-crafted features. Furthermore, we observe that all methods achieve higher accuracy for rooms with fewer bns. These rooms are typically smaller and have only one entrance, resulting in spatially clustered border nodes. In such cases, features considering the border nodes’ semantics lead to improved classification performance. The training accuracy of the GNN is 67.94%, while that of the kNN is 60.50%. Note that the training and validation accuracies reported in Table I correspond to the full set of rooms, irrespective of the number of border nodes.
@@ VI-C Path-Planning on uHumans2 Office Scene
Figure 5: (a) Place subgraph of the uHumans2 office (52m×45m) with its semantic classes (green, blue, red); (b-d) The starting node is denoted by S, and the goal node by G. The path of each algorithm is shown in black. Expanded nodes are depicted in dark gray, while unexpanded nodes in light gray: (b) Path of COA* on the place subgraph; (c) Path of HCOA* on the room subgraph; (d) Path of HCOA* on the place subgraph.
In this section, we perform multiple path-planning scenarios in the office 3DSG shown in Figure 1. The robot must navigate from a start location to a goal while satisfying the following safety constraints, prioritized in descending order: (i) avoid computers, and (ii) avoid entering room R(0), a typically crowded area. To enforce these constraints, we assign semantic labels to place nodes. Specifically, we set $\phi_{V}^{0}(v)=3$ for all $v\in C$, where $C$ denotes the set of place nodes within a radius r=3r=3m of any computer in the scene. Formally, $C=\{v\in G^{0}:\|x(v)-x(o)\|_{2}\leq r,\hskip 5.0pt\forall o\in O\}$ where $x(\cdot)$ denotes the location of a place node $v$ or object $o$, and $O$ is the set of all computers in the 3DSG. We also assign $\phi_{V}^{0}(v)=2$ for all $v\in\text{R(0)}$, thereby penalizing passing through this room. The rest of the nodes have $\phi_{V}^{0}(v)=1$.
First, we evaluate the computational efficiency of the algorithm by performing 1,000 runs with fixed start and goal. The starting node $v_{s}$ is located in room R(2), and the goal $v_{g}$ is in room R(1). Figure 5(5(a)) shows the place layer, where nodes are color-coded according to their semantic class. The computed path of COA* in the place layer is shown in Figure 5(5(b)). Figures 5(5(c))-(5(d)) illustrate the results of HCOA* in the room and place layers. The figures also show the expanded nodes for each algorithm. In particular, COA* expanded nodes in two additional rooms compared to HCOA*.
TABLE II: Path-Planning on uHumans2 Office Scene.
Algorithm Exp. Nodes Time ($10^{-3}$s) Opt. Path
@@
HCOA*-kNN 412 5.6 $\pm$ 0.2
HCOA*-GNN 412 11.5 $\pm$ 2.5
COA* 549 4.9 $\pm$ 0.4
@@
The exact number of expanded nodes is provided in Table II, where HCOA* demonstrates a 25% reduction compared to COA*. Table II also presents the computational time of the algorithms. We observe that HCOA*-MC achieves the best performance in terms of computational efficiency, reducing the execution time by 14% compared to COA*. However, HCOA*-GNN has the highest computational time, as GNN inference can be more time-consuming than graph search in small graphs (the entire place subgraph contains only 1,314 nodes). Additionally, all classification methods classify R(0) as class 2 and R(2) as class 3. However, while GNN and kNN assign class 3 to R(3), MC classifies it as class 1. Although this misclassification does not impact the final path in the current scenario, it could lead to suboptimal solutions in more complex room graphs, as the robustness of the algorithm is highly dependent on the structure of the room graph.
@@
Finally, all four approaches successfully compute the optimal path in the place layer, as shown in Figure 5. It is important to note that while the first assumption of Proposition 2 holds, the second assumption does not. However, this does not affect the results in this case, as the optimal path in the place layer does not require deviations into rooms outside the unique path in the room layer to achieve a lower-cost trajectory.
To further evaluate the suboptimality of HCOA*, we conducted 500 scenarios with varied start and goal locations. HCOA* successfully found the optimal path in 96.95% of the cases. This outcome is highly dependent on the representation of the room layer; modeling a room solely by its centroid can lead to suboptimal paths, particularly in elongated or non-orthogonal rooms (e.g., R(0)) where the centroid may not accurately reflect spatial connectivity (see Figure 3).
We also report results from the MA* implementation. When $\alpha=2$, performance drops to 80.38% (comparing to COA*), while $\alpha=10$ yields 99.84%. For larger values, accuracy approaches 100%, but for $\alpha>10^{5}$, it degrades (e.g., 95.25%), indicating numerical instability. These findings show that MA* requires careful tuning of $\alpha$ based on the graph size, Euclidean distances, and the number of semantic classes, highlighting the benefit of using a total order as in COA* and HCOA*.
VI-D Path-Planning on uHumans2 Subway Scene
Figure 6: (a) uHumans2 subway scene; (b) Place subgraph of the uHumans2 subway (88m×60m) with its semantic classes (green, blue, red); (b-d) The starting node is denoted by S, and the goal node by G. The path of each algorithm is shown in black. Expanded nodes are depicted in dark gray, while unexpanded nodes in light gray: (b) Path of COA* on the place subgraph; (c) Path of HCOA* on the room subgraph; (d) Path of HCOA* on the place subgraph.
We performed similar simulations to those in Section VI-C on the two-floor subway 3DSG, which consists of 2,732 place nodes, shown in Figure 6(6(a)). The safety constraints are prioritized in descending order: (i) avoid room R(5), where stair repairs are in progress, and (ii) avoid room R(0), which is crowded. Accordingly, we assign semantic classes as follows: $\phi_{V}^{0}(v)=3$ for all $v\in\text{R(5)}$, $\phi_{V}^{0}(v)=2$ for all $v\in\text{R(0)}$, and $\phi_{V}^{0}(v)=1$ for all other nodes.
To evaluate the computational efficiency of the algorithms, we performed 1,000 runs with fixed start and goal locations. The starting node $v_{s}$ is located in room R(10), and the goal node $v_{g}$ is in room R(4). Figure 6(6(b)) shows the place layer, while Figures 5(5(b))-(5(d)) depict the computed paths of COA* and HCOA*, along with the corresponding expanded nodes.
Table III summarizes the results. HCOA* achieves a 59% reduction in the number of expanded nodes compared to COA*. In terms of computational time, HCOA*-MC and HCOA*-kNN yield reductions of 49% and 15%, respectively.
We also assessed the suboptimality of the proposed algorithm by running 500 scenarios with varied start and goal locations. HCOA* successfully finds the optimal path in 70.56% of the cases. This decrease compared to the office environment is attributed to the increased complexity introduced by the two-floor layout and the irregular geometry of the rooms. For comparison, the results of MA* show that with $\alpha=2$, the performance is 72.42%, while $\alpha=10$ yields 99.16%. For larger values, accuracy approaches 100%, but degrades for $\alpha>10^{5}$ again indicating numerical instability.
TABLE III: Path-Planning on uHumans2 Subway Scene.
Algorithm Exp. Nodes Time ($10^{-3}$s) Opt. Path
[…]
HCOA*-kNN 1029 15.6 $\pm$ 5.1
HCOA*-GNN 1029 36.6 $\pm$ 4.4
COA* 2480 18.3 $\pm$ 4.5
VII Conclusion